# Dynamic Conditional Optimal Transport through Simulation-Free Flows

### v1

@@ 3.1 Unconditional Optimal Transport
We provide here a brief overview of optimal transport in the standard unconditional setting. For more details, we refer to the standard references of Villani et al. (2009); Santambrogio (2015); Ambrosio et al. (2005). Fix a cost function $c:X\times X\to\mathbb{R}\cup\{+\infty\}$. Suppose we have two Borel measures $\eta,\nu\in\mathbb{P}(X)$. The Monge Problem seeks to find a measurable transport map $T:X\to X$ minimizing the expected cost of transport, i.e. corresponding to the optimization problem
@@
Here, we constrain our minimization problem over the set of measures and vector fields $(\gamma_{t},v_{t})$ interpolating between $\eta$ and $\nu$, satisfying a continuity equation (see Section 5.1). In Section 4, we study a generalization of the Wasserstein distances for conditional optimal transport problems. In particular, Theorem 16 provides a generalization of the Benamou-Brenier theorem to the conditional setting which recovers a conditional Wasserstein distance.
3.2 Static Conditional Optimal Transport
In the unconditional setting, one seeks an optimal coupling between two given measures. In contrast, in the conditional setting, we would like to find a family of couplings that can transport a given source measure into any conditional measure of the target distribution. In other words, given a target measure $\nu\in\mathbb{P}(Y\times U)$ and some source measure $\eta\in\mathbb{P}(U)$, one seeks a transport map $T:Y\times U\to U$ such that
$T_{\#}(y,{-})_{\#}\eta=\nu^{y}\qquad\forall y\in Y.$ (5)
If such a map were available, by drawing samples $u_{0}\sim\eta$ and transforming them, one would obtain samples $T(y,u)\sim\nu^{y}$. Solving this transport problem for each fixed $y$ is expensive at best, or impossible when only has a single (or no) samples $(y,u)\sim\nu$ for any given $y$.
Thus, one must leverage information across different observations $y$. To that end, recent work has focused on the notion of triangular mappings $T:Y\times U\to Y\times U$ (Hosseini et al., 2023; Baptista et al., 2020) (Bogachev and Ruas, 2007, Section 10.10) of the form
$T(y,u)=\left(T_{Y}(y),T_{U}(T_{Y}(y),u)\right)$ (6)
for some $T_{Y}:Y\to Y$ and $T_{U}:Y\times U\to U$. Triangular mappings are of interest as they allow us to obtain conditional couplings from joint couplings.
Proposition 1 (Theorem 2.4 (Baptista et al., 2020), Prop. 2.3 (Hosseini et al., 2023)).
Definition 2 (Conditional Wasserstein Space).
Suppose $\mu\in\mathbb{P}(Y)$ is a given probability measure over $Y$ and $1\leq p<\infty$. The set
$\mathbb{P}^{\mu}(Y\times U)=\left\{\gamma\in\mathbb{P}_{p}(Y\times U)\mid\pi^{Y}_{\#}\gamma=\mu\right\}$ (7)
is the space of joint measures on $Y\times U$ having fixed $Y$-marginals $\mu$. The set $\mathbb{P}_{p}^{\mu}(Y\times U):=\mathbb{P}^{\mu}(Y\times U)\cap\mathbb{P}_{p}(Y\times U)$ is the conditional $p$-Wasserstein space, consisting of joint measures with finite $p$th moments and $Y$-marginals $\mu$.
Given a source and target measures $\eta,\nu\in\mathbb{P}^{\mu}(Y\times U)$, the conditional Monge problem seeks to find a triangular mapping solving
@@ 3.2 Static Conditional Optimal Transport
and we define the set of $(Y)$-restricted probability measures $\mathcal{R}_{Y}\subset\mathbb{P}\left((Y\times U)^{2}\right)$ such that every $\gamma\in\mathcal{R}_{Y}$ is concentrated on $\mathscr{C}$. In other words, if $\gamma\in\mathcal{R}_{Y}$, then samples $(y_{0},u_{0},y_{1},u_{1})\sim\gamma$ have $y_{0}=y_{1}$ almost surely. In addition, for any $\eta,\nu\in\mathbb{P}(Y\times U)$, we define the set of $(Y)$-restricted couplings $\Pi_{Y}(\eta,\nu)$ to be the probability measures in $\mathcal{R}_{Y}$ whose marginals are $\eta$ and $\nu$, i.e.
@@
Hosseini et al. (2023) prove the existence of minimizers to the conditional Kantorovich and Monge problems under very general assumptions. Moreover, optimal couplings to the conditional Kantorovich problem induce optimal couplings for $\mu$-almost every conditional measure. Assuming sufficient regularity assumptions on the conditional measures, unique solutions to the conditional Monge problem exist.
Proposition 3 (Prop 3.3 (Hosseini et al., 2023)).
@@
Proposition 4 (Prop 3.8 (Hosseini et al., 2023)).
@@
4 Conditional Wasserstein Space
We now introduce a metric on the space $\mathbb{P}_{p}^{\mu}(Y\times U)$. Intuitively, the conditional Wasserstein distance measures the usual Wasserstein distance between all of the conditional distributions in expectation under the fixed $Y$-marginal $\mu$.
Definition 5 (Conditional $p$-Wasserstein Distance).
Suppose $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ and $1\leq p<\infty$. The function
is the conditional $p$-Wasserstein distance. Here, $W_{p}^{p}:\mathbb{P}_{p}(U)\times\mathbb{P}_{p}(U)\to\mathbb{R}$ is the usual $p$-Wasserstein distance for measures on $U$.
By an application of Jensen’s inequality we observe that $W_{p}^{\mu}(\eta,\nu)\geq\mathbb{E}_{y\sim\mu}\left[W_{p}(\eta^{y},\nu^{y})\right]$. For $p>1$, this inequality is strict unless $W_{p}(\eta^{y},\nu^{y})$ is $\mu$-almost surely constant. We emphasize that Definition 5 is purely formal at this stage. We do not know if this integral is well-defined, or if it actually defines a metric. Before proceeding, we first note that $W_{p}^{\mu}(\eta,\nu)$ may be viewed as the minimal value of the constrained Kantorovich problem in Equation (11) when one takes the cost to be the metric on the space $Y\times U$. Although Hosseini et al. (2023) studied the existence of minimizers to this problem, they did not study the metric properties of the resulting space.
Proposition 6 (Equivalent Formulation of the $p$-Wasserstein Distance).
@@ 4 Conditional Wasserstein Space
We now show that $W_{p}^{\mu}$ is a metric on $\mathbb{P}_{p}^{\mu}(Y\times U)$.
Proposition 7 ($W_{p}^{\mu}$ is a Metric).
For $1\leq p<\infty$, the conditional $p$-Wasserstein distance $W_{p}^{\mu}$ is a metric on $\mathbb{P}_{p}^{\mu}(Y\times U)$.
@@
Fix $\eta,\nu,\rho\in\mathbb{P}_{p}^{\mu}(Y\times U)$. Since $W_{p}$ is a metric on $\mathbb{P}_{p}(U)$, we immediately obtain the symmetry of $W_{p}^{\mu}$. Moreover, we have that $W_{p}^{\mu}(\eta,\nu)=0$ if and only if $\eta^{y}=\nu^{y}$ for $\mu$-almost every $y$. Thus, if $W_{p}^{\mu}(\eta,\nu)=0$ and $E\subseteq Y\times U$ is Borel measurable,
@@
4.1 Metric Properties of Conditional Wasserstein Distances
We now briefly explore some properties of the space $\mathbb{P}_{p}^{\mu}(Y\times U)$ equipped with the conditional Wasserstein distance $W_{p}^{\mu}$. In particular, we observe that the topology on $\mathbb{P}_{p}^{\mu}(Y\times U)$ generated by $W_{p}^{\mu}$ is stronger than that generated by $W_{p}$. Note that we easily see $W_{p}^{\mu}(\eta,\nu)\geq W_{p}(\eta,\nu)$ for every $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ by the characterization in Proposition 6. The converse is, however, false, and the metrics are inequivalent.
Proposition 8 ($W_{p}^{\mu}$ is stronger than $W_{p}$).
There does not in general exist a constant $C>0$ such that $W_{p}^{\mu}(\eta,\nu)\leq CW_{p}(\eta,\nu)$ for all $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$.
Proof.
We provide a counterexample. Fix any $u_{0}\neq 0\in U$ and $y_{0},y_{1}\in Y$ such that $y_{0}\neq y_{1}$. Define $\mu=\frac{1}{2}\left(\delta_{y_{0}}+\delta_{y_{1}}\right)$. Set $u_{k}=(k+1)u_{0}$ for $k=1,2,\dots$ and for each $k$, define two measures on $Y\times U$ by
@@ 4.1 Metric Properties of Conditional Wasserstein Distances
Moreover, as $k\to\infty$ we have $W_{p}^{\mu}(\mu_{k},\nu_{k})\to\infty$ but $W_{p}^{p}(\nu_{k},\eta_{k})$ remains bounded. ∎
In addition, we can show that the conditional Wasserstein distance dominates the Wasserstein distance on the $U$ marginals.
Proposition 9 (Conditional Wasserstein Dominates $U$-Marginal Distance).
For $1\leq p<\infty$ and $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, we have
$W_{p}\left(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu\right)\leq W_{p}^{\mu}(\eta,\nu).$ (25)
Proof.
@@
since $\gamma\in\Pi(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ is a coupling but potentially sub-optimal. ∎
Example: Gaussian Measures.
We conclude with an example where the conditional $2$-Wasserstein distance may be explicitly computed. In particular, this example shows that the inequality in Proposition 9 may be strict. Suppose $Y$ and $U$ are Euclidean spaces (of possibly different dimensions), and that $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ are Gaussians of the form
It follows that $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu=\mathcal{N}(m,\Sigma)$. Because we may obtain $\eta^{y},\nu^{y}$ in closed form, we can compute $W_{p}^{\mu}(\eta,\nu)$ in closed form. Let
$Q^{\eta}=\Sigma_{22}^{\eta}-\Sigma_{21}^{\eta}\Sigma^{-1}\Sigma_{12}^{\eta}\qquad Q^{\nu}=\Sigma_{22}^{\nu}-\Sigma_{21}^{\nu}\Sigma^{-1}\Sigma_{12}^{\nu}\qquad R=(\Sigma_{21}^{\eta}-\Sigma_{21}^{\nu})\Sigma^{-1}$ (34)
Then, we obtain by a brute-force calculation that
$W_{2}^{\mu,2}(\eta,\nu)=|m_{u}^{\eta}-m_{u}^{\nu}|^{2}+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}+R\Sigma R^{T}\right).$ (35)
Note that when $\eta,\nu$ have uncorrelated $Y,U$ components, we precisely recover $W_{2}^{2}(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ as one may expect. As a special case of interest, if $Y=U=\mathbb{R}$ and
then we obtain as a special case of Equation (35) that $W_{2}^{\mu,2}(\eta,\nu)=2(1-\sqrt{1-\rho^{2}})$. This is zero if and only if $\rho=0$, i.e. $\eta=\nu$. However, $\pi^{U}_{\#}\eta=\pi^{U}_{\#}\nu=\mathcal{N}(0,1)$.
4.2 Conditional Wasserstein Space as a Geodesic Space
In this section, we show that $\mathbb{P}_{p}^{\mu}(Y\times U)$ is a geodesic space. In fact, we show the stronger statement that there exists a constant speed geodesic between any two measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$, which is a generalization of known results (Santambrogio, 2015, Theorem 5.27). First, we introduce some preliminary notions. A curve is a continuous function $\gamma_{\bullet}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ where $I=(a,b)\subseteq\mathbb{R}$ is any open interval of finite length. A curve is said to be absolutely continuous if there exists $m\in L^{1}((a,b))$ such that
$W_{p}^{\mu}(\gamma_{s},\gamma_{t})\leq\int_{s}^{t}m(\tau)\,\mathrm{d}\tau\qquad\forall a<s\leq t<b.$ (37)
If $(\gamma_{t})$ is an absolutely continuous curve, then its metric derivative
$|\gamma^{\prime}|(t)=\lim_{s\to t}\frac{W_{p}^{\mu}(\gamma_{s},\gamma_{t})}{|s-t|}$ (38)
exists for almost every $t\in(a,b)$, and, moreover, we almost surely have $|\gamma^{\prime}|(t)\leq m(t)$ pointwise for any $m$ satisfying Equation (37) (Ambrosio et al., 2005, Theorem 1.1.2). A curve $(\gamma_{t})$ is called a constant speed geodesic if for all $a<s\leq t<b$, we have
$W_{p}^{\mu}(\gamma_{s},\gamma_{t})=|t-s|W_{p}^{\mu}(\gamma_{a},\gamma_{b}).$ (39)
It is straightforward to show that every constant speed geodesic is absolutely continuous.
Theorem 10 ($\mathbb{P}_{p}^{\mu}(Y\times U$) is a Geodesic Space).
$\mathbb{P}_{p}^{\mu}(Y\times U)$ admits constant-speed geodesics. That is, for any $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, there exists a constant speed geodesic, contained within $\mathbb{P}_{p}^{\mu}(Y\times U)$, between $\eta$ and $\nu$.
@@ 4.2 Conditional Wasserstein Space as a Geodesic Space
When an optimal restricted coupling $\gamma^{\star}\in\Pi_{Y}(\eta,\nu)$ is induced by an injective triangular map $T^{\star}$, we may recover a constant speed geodesic in $\mathbb{P}_{p}^{\mu}(Y\times U)$, generalizing the McCann interpolant (McCann, 1997) to the conditional setting. We refer to Proposition 4 for sufficient conditions on $\eta,\nu$ under which such a $T^{\star}$ exists. Informally, samples from $(y_{0},u_{0})\sim\eta$ flow in a straight path at a constant speed to their destination $T^{\star}(y_{0},u_{0})$. We note that this theorem, while primarily about geodesics, relies on notions discussed in Section 5. In fact, this may be viewed as a special case of the results in Section 5, where we may explicitly identify the vector field.
Theorem 11 (Conditional McCann Interpolants).
Fix $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$. Suppose $T^{\star}(y,u)=(y,T^{\star}_{\mathcal{U}}(y,u))$ is a block triangular map solving the conditional Monge problem (8). Define the maps $T_{t}:Y\times U\to Y\times U$ for $0\leq t\leq 1$ via
$T_{t}=(1-t)I+tT^{\star}\qquad(y,u)\mapsto(y,(1-t)u+tT^{\star}_{\mathcal{U}}(y,u))$ (49)
and define the curve of measures $\gamma_{t}=[T_{t}]_{\#}\eta\in\mathbb{P}_{p}^{\gamma}(Y\times U)$. Then,
  1. 1.
  2. 2.
@@
5 Conditional Benamou-Brenier Theorem
In this section, we prove a characterization of all absolutely continuous curves in the space $\mathbb{P}_{p}^{\mu}(Y\times U)$. Here we assume $1<p<\infty$. First, we introduce the notion of a triangular vector field on $Y\times U$, which informally has no velocity in the $Y$ component. Triangular vector fields serve as the dynamic analog of triangular maps (6).
Definition 12 (Triangular Vector Fields).
Let $I\subset\mathbb{R}$ be an open interval. A time-dependent Borel vector field $v:I\times Y\times U\to Y\times U$ is said to be triangular if there exists a Borel vector field $v^{U}:I\times Y\times U\to U$ such that
$v_{t}(y,u)=\left(0,v_{t}^{U}(y,u)\right).$ (60)
In other words, $v_{t}=\iota\left(v_{t}^{U}\right)$ where $\iota:U\to Y\times U$ is the inclusion map.
5.1 Continuity Equation
Let $I\subset\mathbb{R}$ be an open interval. The continuity equation
$\partial_{t}\gamma_{t}+\text{div}(v_{t}\gamma_{t})=0\qquad\text{on }Y\times U\times I$ (61)
describes the evolution of a measure $\gamma_{t}$ which flows along a given vector field $v_{t}$. This equation must be understood distributionally, i.e. for every $\varphi$ in an appropriate space of test functions,
We now prove a lemma that will be critical to our later results. Informally, a solution to the continuity equation with a triangular vector field will result in the conditional measures almost surely satisfying the continuity equaiton as well.
Lemma 13 (Triangular Vector Fields Preserve Conditionals).
Suppose $v_{t}(y,u)=(0,v_{t}^{U}(y,u))$ is triangular and that $(\gamma_{t})\subset\mathbb{P}_{p}^{\mu}(Y\times U)$ is a path of measures such that $(v_{t},\gamma_{t})$ satisfy the continuity equation (61) in the distributional sense. Then, it follows that for $\mu$-almost every $y\in Y$, we have
$\partial_{t}\gamma_{t}^{y}+\nabla\cdot(v_{t}^{U}(y,{-})\gamma_{t}^{y})=0.$ (63)

### v2

@@ Appendix A Unconditional Optimal Transport
We provide here a brief and informal overview of optimal transport in the standard unconditional setting. For more details, we refer to the standard references of Villani et al. (2009); Santambrogio (2015); Ambrosio et al. (2005). Let $X$ be a separable metric space and fix a cost function $c:X\times X\to\mathbb{R}\cup\{+\infty\}$. Suppose we have two Borel measures $\eta,\nu\in\mathbb{P}(X)$. The Monge problem seeks to find a measurable transport map $T:X\to X$ minimizing the expected cost of transport, i.e. corresponding to the optimization problem
@@
Here, we constrain our minimization problem over the set of measures and vector fields $(\gamma_{t},v_{t})$ interpolating between $\eta$ and $\nu$, satisfying a continuity equation (see Section 5). In Section 4, we study a generalization of the Wasserstein distances for conditional optimal transport problems. In particular, Theorem 5 provides a generalization of the Benamou-Brenier theorem to the conditional setting which recovers a conditional Wasserstein distance.
Appendix B Conditional Optimal Transport
This section contains additional discussion regarding the static COT problem, supplementing Section 3.1. We refer to Hosseini et al. (2023), Baptista et al. (2020), and Chemseddine et al. (2023) for further results and details.
Given a source and target measures $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, the conditional Monge problem seeks to find a triangular mapping solving
@@ Appendix B Conditional Optimal Transport
and we define the set of $(Y)$-restricted probability measures $\mathcal{R}_{Y}\subset\mathbb{P}\left((Y\times U)^{2}\right)$ such that every $\gamma\in\mathcal{R}_{Y}$ is concentrated on $\mathscr{C}$. In other words, if $\gamma\in\mathcal{R}_{Y}$, then samples $(y_{0},u_{0},y_{1},u_{1})\sim\gamma$ have $y_{0}=y_{1}$ almost surely. In addition, for any $\eta,\nu\in\mathbb{P}(Y\times U)$, we define the set of triangular couplings $\Pi_{Y}(\eta,\nu)$ to be the probability measures in $\mathcal{R}_{Y}$ whose marginals are $\eta$ and $\nu$, i.e.
@@
Hosseini et al. (2023) prove the existence of minimizers to the conditional Kantorovich and Monge problems under very general assumptions. Moreover, optimal couplings to the conditional Kantorovich problem induce optimal couplings for $\mu$-almost every conditional measure. Assuming sufficient regularity assumptions on the conditional measures, unique solutions to the conditional Monge problem exist. We restate these results here for the sake of completeness.
Proposition 3 (Prop 3.3 (Hosseini et al., 2023))
@@
Proposition 4 (Prop 3.8 (Hosseini et al., 2023))
@@
Appendix C Closed-Form Conditional Wasserstein Distance for Gaussian Measures
In this section, we provide additional details and results regarding the closed-form conditional Wasserstein distance for Gaussian distributions. See Section 4 in the main paper.
Suppose $Y$ and $U$ are Euclidean spaces (of possibly different dimensions), and that $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ are Gaussians of the form
This form is chosen to ensure that $\eta$ and $\nu$ have equal $Y$-marginals. It follows that $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu=\mathcal{N}(m,\Sigma)$. Let
$Q^{\eta}=\Sigma_{22}^{\eta}-\Sigma_{21}^{\eta}\Sigma^{-1}\Sigma_{12}^{\eta}\qquad Q^{\nu}=\Sigma_{22}^{\nu}-\Sigma_{21}^{\nu}\Sigma^{-1}\Sigma_{12}^{\nu}\qquad R=(\Sigma_{21}^{\eta}-\Sigma_{21}^{\nu})\Sigma^{-1}.$ (21)
We have that the conditionals $\eta^{y},\nu^{y}$ are available in closed-form:
$\eta^{y}=\mathcal{N}\left(m_{u}^{\eta}+\Sigma_{21}^{\eta}\Sigma^{-1}(y-m),Q^{\eta}\right)\qquad\nu^{y}=\mathcal{N}\left(m_{u}^{\nu}+\Sigma_{21}^{\nu}\Sigma^{-1}(y-m),Q^{\nu}\right).$ (22)
Thus, for any fixed $y$, we use the known closed-form unconditional Wasserstein distance to obtain
$\begin{split}W_{2}^{2}(\eta^{y},\nu^{y})=&\big{|}m_{u}^{\eta}-m_{u}^{\nu}+R(y-m)\big{|}^{2}+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}\right).\end{split}$ (23)
We now take an expectation over $y\sim\mu=\mathcal{N}(m,\Sigma)$ to compute $W_{2}^{\mu,2}$. Observe that $R(y-m)\sim\mathcal{N}(0,R\Sigma R^{\sf T})$ and that $\mathbb{E}_{y\sim\mu}[|R(y-m)|^{2}]=\text{Tr}(R\Sigma R^{\sf T})$. Thus,
$W_{2}^{\mu,2}(\eta,\nu)$ $=\mathbb{E}_{y\sim\mu}\left[W_{2}^{2}(\eta^{y},\nu^{y})\right]$ (24)
$=\mathbb{E}_{y\sim\mu}\left[|m_{u}^{\eta}-m_{u}^{\nu}|^{2}+2\langle m_{u}^{\eta}-m_{u}^{\nu},R(y-m)\rangle+|R(y-m)|^{2}\right]$ (25)
$\qquad+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}\right)$
$=|m_{u}^{\eta}-m_{u}^{\nu}|^{2}+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}+R\Sigma R^{T}\right).$ (26)
This form, perhaps unsurprisingly, closely resembles the unconditional Wasserstein distance between two Gaussians, except for the presence of an additional $\text{Tr}(R\Sigma R^{\sf T})$ term. Note that when $\eta,\nu$ have uncorrelated $Y,U$ components, we precisely recover $W_{2}^{2}(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ as one may expect. As a special case of interest, if $Y=U=\mathbb{R}$ and
then we obtain as a special case of Equation (24) that $W_{2}^{\mu,2}(\eta,\nu)=2(1-\sqrt{1-\rho^{2}})$. This is zero if and only if $\rho=0$, i.e. $\eta=\nu$.
Appendix D Proofs: Section 4
In this section, we provide detailed proofs of our claims in Section 4, regarding the metric properties of the conditional Wasserstein space.
D.1 Metric Properties
We first note that $W_{p}^{\mu}(\eta,\nu)$ may be viewed as the minimal value of the constrained Kantorovich problem in Equation (3) when one takes the cost to be the metric on the space $Y\times U$. Similar results, relating the conditional Wasserstein distance to triangular couplings, have appeared previously, but our proof is independent of these prior works (Chemseddine et al., 2023; Gigli, 2008).
Proposition 5 (Equivalent Formulation of the Conditional Wasserstein Distance)
@@ D.1 Metric Properties
We now proceed to prove several metric properties of our distance.
See 2
@@
Part (a). This is simply a restatement of Proposition 5.
Part (b). Fix $\eta,\nu,\rho\in\mathbb{P}_{p}^{\mu}(Y\times U)$. Since $W_{p}$ is a metric on $\mathbb{P}_{p}(U)$, we immediately obtain the symmetry of $W_{p}^{\mu}$. Moreover, we have that $W_{p}^{\mu}(\eta,\nu)=0$ if and only if $\eta^{y}=\nu^{y}$ for $\mu$-almost every $y$. Thus, if $W_{p}^{\mu}(\eta,\nu)=0$ and $E\subseteq Y\times U$ is Borel measurable,
@@
Part (c). We provide a counterexample. Fix any $u_{0}\neq 0\in U$ and $y_{0},y_{1}\in Y$ such that $y_{0}\neq y_{1}$. Define $\mu=\frac{1}{2}\left(\delta_{y_{0}}+\delta_{y_{1}}\right)$. Set $u_{k}=(k+1)u_{0}$ for $k=1,2,\dots$ and for each $k$, define two measures on $Y\times U$ by
@@
Moreover, as $k\to\infty$ we have $W_{p}^{\mu}(\mu_{k},\nu_{k})\to\infty$ but $W_{p}^{p}(\nu_{k},\eta_{k})$ remains bounded. See Figure 4.
Part (d). First, the unconditional distance $W_{p}(\eta,\nu)$ may be obtained via an unrestricted coupling in $\Pi(\eta,\nu)$, i.e. the set of all joint measures on $Y\times U$ having marginals $\eta,\nu$. Since $\Pi(\eta,\nu)\supseteq\Pi_{Y}(\eta,\nu)$, by part (a) we see that $W_{p}(\eta,\nu)\leq W_{p}^{\mu}(\eta,\nu)$.
@@
since $\gamma\in\Pi(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ is a coupling but potentially sub-optimal.
$Y$$U$$y_{0}$$y_{1}$$u_{0}$$u_{k}$
Figure 4: The counterexample in Proposition 2. The measure $\eta_{k}$ is shown in black and the measure $\nu_{k}$ is shown in white.
D.2 Geodesics
We now study the geodesics in the space $\mathbb{P}_{p}^{\mu}(Y\times U)$.
See 1
@@ D.2 Geodesics
See 2
@@
Appendix E Proofs: Section 5
In this section, we provide proofs of all claims made in Section 5.
E.1 Continuity Equation
We begin with a lemma that is used in the proof of Theorem 4. Informally, a solution to the continuity equation with a triangular vector field will result in the conditional measures almost surely satisfying the continuity equation as well.
See 1